OpenStudy (anonymous):

how to solve Ln sqr x/2x+3

OpenStudy (anonymous):

\[\ln(\sqrt{\frac{x}{x+3}})\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what are you supposed to do with it?